OpenStudy (anonymous):

need help here is someone can help me with my math A person hired a firm to build a CB tower the firm charges 100$ for labor for the first 10 feet, after that the cost of the labor for each succeeding 10 feet is 25more than the preceding 10 feet , that is the next 10 feet will cost $125,the next 10 feet will cost 150,etc.how much will it cost to build a 90 feet tower? need help please

OpenStudy (anonymous):

100+25+25+125+150

OpenStudy (anonymous):

this is my first math in 15 years just need help if anyone can help me with this

OpenStudy (anonymous):

10 feet charge = $100 20 feet charge = $125 30 feet charge = $150 40 feet charge = $175 50 feet charge = $200 60 feet charge = $225 70 feet charge = $250 80 feet charge = $275 90 feet charge = $300 The cost to build a 90-foot tower is the summation of all above. That is =100+125+150+175+200+225+250+275+300 =$ 1800 The cost will be $1800

OpenStudy (anonymous):

all i did was copy and paste question in a google searcH

OpenStudy (anonymous):

are you sure this the right answer ? and that's how it was explained .. is there a formula?

OpenStudy (anonymous):

This problem involves arithmetic sequence since labor cost for each successive 10 ft remains constant at $25 The arithmetic sequence of the labor costs is 100, 125, 150, … For finding the cost for building a 90 ft tower, we sum the first 9 terms of the above sequence a = 100, d = 25, n = 9 Sn = (n/2)[2a + (n - 1)d] S9 = (9/2)[2 * 100 + (9 - 1)(25)] = 1800 The cost will be $1800 Source: ww2.justanswer.com/.../2010-11-15_071033...

OpenStudy (anonymous):

man you the best ... Thanks brother can it help me with other questions ..